4. Suppose that of 80 students who studied more, 50 of them got better grades, and of the 60 students who studied less, 35 got better grades. Fill in the table below and calculate the rate difference and relative risk. Use your results to assess the risk posed by studying. Is it possible that students who study more are at risk of better grades, based on these data?

Better Grades

Worse Grades

Total

Study more

Study less

Total
Answer #1
Data interpretation
Better grades Worse grades Total
Study more 50 30 80
Study less 35 25 60
Total 85 55 140

So. there are 80 students who study more among them 50 students have a better grade and 30 students have worse grade.

60 students who studied less among them 35 got a better grade and 25 have less grade.

Here rate difference will be

The incidence rate among exposed(Study more) - The incidence rate among no exposed(Study less)

50:150 + 30 = 0.625 is the incidence rate of exposed

35:35 +25) = 0.583 is the incidence rate of no exposed

So Rate difference will be

  0.625 -0.583 = 0.042

Relative risk

The incidence rate among exposed(Study more)/The incidence rate among no exposed(Study less)

0.625 -0.583 = 1.07

Yes, there is a 7% higher risk of better grades among them who studies more.